The intensity of the light from an object varies inversely with the square of the distance. Two stars give off the same amount of light. However from earth their intensities differ. Let I1 and I2 be there intensities and let d1 and d2 be their respective distances from earth. What is the ratio of d2 to d1?
Answer by TimothyLamb(4379)

i = C/dd
dd = C/i
d = sqrt(C/i)
d = sqrt(C) * sqrt(1/i)
---
where C is a constant of proportionality
---
d1 = sqrt(C) * sqrt(1/I1)
d2 = sqrt(C) * sqrt(1/I2)
---
d2/d1 = ( sqrt(C) * sqrt(1/I2) )/( sqrt(C) * sqrt(1/I1) )
d2/d1 = ( sqrt(C)/(sqrt(C) ) * ( sqrt(1/I2)/sqrt(1/I1) )
d2/d1 = 1 * ( sqrt(1/I2)/sqrt(1/I1) )
---
answer:
d2/d1 = sqrt(1/I2)/sqrt(1/I1)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php